Calculate the grams of sulfur in 8.022x1025 formula units of Lithium sulfate.


Please answer with correct number of sig figs, and scientific notation


1
Expert's answer
2021-02-10T02:30:28-0500

A mole of Lithium sulfate contains 6.02×1023 molecules.

Proportion 1:

1 mol - 6.02×1023 molecules

x mol - 8.022×1025 molecules

"x = \frac{1 \times 8.022 \times 10^{25}}{6.02 \times 10^{23}} \\

= 133.25 \;mol"

M(Li2SO4) = 109.94 g/mol

m(Li2SO4) "= n \times M"

"= 133.25 \times 109.94 \\

= 14649.50 \;g"

M(S) = 32.06 g/mol

Proportion 2:

109.94 – 100 %

32.06 – y

"y = \frac{32.06 \times 100}{109.94} = 29.16 \%"

Proportion 3:

14649.50 – 100 %

z – 29.16 %

"z = \frac{14649.50 \times 29.16}{100}=4271.79 \;g"

Answer: 4271.79 g